OpenStudy (pottersheep):

Grade 11 Sequences help pleasee!

OpenStudy (pottersheep):

p-4, 8p + 3 , 10p - 5 are the 4th, 6th and 8th terms of an arithmetic sequence. Find: a) 1st term of the sequence b) common difference

OpenStudy (anonymous):

solve the equation p-4=a+3d 8p+3=a+5d 10p-5=a+7d a is the first term and d is the common difference

OpenStudy (pottersheep):

I do do those things but that answer for a) never matches the right answer in my book. The answer for part a) is 14.

OpenStudy (pottersheep):

I get weird fraction numbers...

OpenStudy (anonymous):

p=-3, d=-7, a=14, what is your answer?

OpenStudy (pottersheep):

for d I got 7(p+1) /2 a I got 19p + 29/2

OpenStudy (anonymous):

there are three equations, plug a and d in them, and then you work out p

OpenStudy (pottersheep):

so then to find p I did p - 4 (which is term 4) = a + 3d p - 4 = (19 + 29) /2 + 3(7p+7)/2 2p-8 = 19p + 29 + 21p + 21 2p - 8 = 40 p + 50............. -38 p = 58 :(

OpenStudy (anonymous):

hey i found the error! a=- (19p+29)/2

OpenStudy (pottersheep):

:O....

OpenStudy (pottersheep):

Thank you !! I wish I could give you more medals !! I spent so much time on thiss ahhhhh at least it was a rather minor error........meh THANKS!!

OpenStudy (anonymous):

haha!